On June 30, 2013, Sarasota Limited issued 11.25% bonds with a par value of $754,000 due in 20 years. They were issued at 96 and were callable at 102 at any date after June 30, 2020.

Because of lower interest rates and a significant change in the company’s credit rating, it was decided to call the entire issue on June 30, 2020, and to issue new bonds. New 10% bonds were sold in the amount of $1 million at 103; they mature in 20 years. The company follows ASPE and uses straight-line amortization. The interest payment dates are December 31 and June 30 of each year.

(a)

Prepare journal entries to record the retirement of the old issue and the sale of the new issue on June 30, 2020. (Credit account titles are automatically indented when the amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts.)

Date

Account Titles and Explanation

Debit

Credit

June 30, 2020

enter an account title to record redemption of bonds payable

enter a debit amount

enter a credit amount

enter an account title to record redemption of bonds payable

enter a debit amount

enter a credit amount

enter an account title to record redemption of bonds payable

enter a debit amount

enter a credit amount

(To record redemption of bonds payable)

June 30, 2020

enter an account title to record issuance of new bonds

enter a debit amount

enter a credit amount

enter an account title to record issuance of new bonds

enter a debit amount

enter a credit amount

(To record issuance of new bonds)
Answer #1

Answer:

Date Particulars Debit ($) Credit ($)
June 30 2020 Bonds payable 754,000
Loss on redemption of bonds 34684
Discount on bonds payable 754000*0.04 = 30160 - (30160/20*7) = 10556
19604
Cash (754000*102%) 769,080
(To record redemption of bonds payable)
June 30 2020 Cash (1000000*103%) 1,030,000
Premium on bonds payable 30,000
Bonds payable 1,000,000
(To record issuance of new bonds)